Solve. Be sure to show the equations used for the solution.
Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?
Let x = nickels
Let y = dimes

Show your solution below:

How many nickels?
How many dimes?

Answer by stanbon(75887)

Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?
Let x = nickels
Let y = dimes
-----------------------
EQUATIONS:
Quantity: x+y = 50
Value : 5x + 10y = 350
----------------------------
y = 50-x
Substitute:
5x + 10(50-x) = 350
5x + 500 - 10x = 350
-5x = - 150
x = 30 (# of nickels)
y = 50-x = 20 (# of dimes)
-------------------------------
Cheers,
Stan H.